According to one pollster, 53% of children are afraid of the dark. Suppose that a sample of size 30 is drawn. Find the value of  , the standard deviation of the distribution of sample proportions.

Write only a number as your answer. Round to two decimal places (for example: 8.21).
Answer #1

Solution

Given that,

p = 53%=0.53

1 - p = 1-0.53=0.47

n = 30

= p =0.53

standard deviation of the distribution of sample proportion

=  [p ( 1 - p ) / n] =   [(0.53*0.47) /30 ] = 0.09